A positive charge of 4.80 μC is fixed in place. From a distance of 3.40 cm a particle of mass 5.20 g and charge +3.70 μC is fired with an initial speed of 66.0 m/s directly toward the fixed charge. How close to the fixed charge does the particle get before it comes to rest and starts traveling away?
Answer #1

Solution :-

Given data:-

first charge = 4.8 uC

second charge = 3.7 uC

distance = 4.4 cm

mass = 5.2 gm

velocity = 66 m/s

initial energy = kinetic energy + electric potential energy

initial energy = 0.5 * 5.2 * 10^-3 * 66^2 + 9 * 10^9 * 4.8* 10^-6 * 3.7 * 10^-6 / 0.034

initial energy = 16.026 J

final energy = electric potential energy

final energy = 9 * 10^9 * 4.8* 10^-6 * 3.7 * 10^-6 / x

by conservation of energy

final energy = initial energy

9 * 10^9 * 4.8 * 10^-6 * 3.7 * 10^-6 / x = 16.026

x = 0.00997m

particle will come 0.997 cm close to fixed